# AI-Driven Guided Response for Security Operation Centers with Microsoft Copilot for Security

### v3

@@ 5.2. Triage Recommendations
In the triage phase, we leverage incident embeddings produced during preprocessing, along with the latest version of the triage model, to generate triage recommendations. Each incident is evaluated by the model and given a prediction of true positive (TP), false positive (FP), or benign positive (BP), the latter considered as an informational incident. The confidence of each recommendation is assessed against a precision threshold of to ensure that only reliable recommendations are sent to SOC analysts.
@@ 5.3. Investigation Recommendations
In the investigation phase, we utilize the incident embeddings produced during the preprocessing step to generate recommendations for similar incidents. This process begins with the retrieval of historical incident embeddings from our Azure Data Lake Storage (ADLS), going back up to 180 days. These embeddings capture past incidents in a vectorized format, enabling efficient comparison. The core of our approach is to match new incidents with historically relevant incidents within the same organization through a three-step matching process:
@@
Top-k similar incident selection. We select the top-k most similar incidents, up to a maximum of five. Exact and cosine similarity matches are ordered, with a higher priority given to exact matches to ensure the most germane comparisons.
@@ 5.4. Remediation Recommendations
In the remediation phase, we leverage alert embeddings produced during preprocessing, along with the latest version of the remediation model to generate targeted response action recommendations. Each alert is evaluated by the model and given a prediction of ‘contain user’, ‘isolate machine’, or ’stop virtual machine’. The confidence of each recommendation is assessed against a precision threshold of to ensure that only reliable recommendations are provided to the SOC analyst. Next, the system identifies the correct entity (e.g., user, device, virtual machine) associated with each alert based on a set of encoded rules that leverage security research domain knowledge. Finally, individual alert recommendations are aggregated to create comprehensive incident recommendations.
@@ 6. Embedding Pipeline
This pipeline generates historical incident embeddings that allow the similar incident recommendation algorithm to leverage up to 180 days of historical data when making recommendations. Due to the limitations of the training pipeline in processing huge volumes of historical incident telemetry across large regions, we developed a specialized mechanism to generate historical embeddings. This approach guarantees that our similar incident recommendation algorithm rapidly reaches comprehensive historical incident coverage each time the training pipeline is executed.
Continuous embedding generation. The embedding pipeline operates in a continuous loop, with each iteration processing data from one day further back than the last. Leveraging the preprocessing steps outlined in Section 5.1, we integrate a deduplication process. This involves loading historical incident embeddings and comparing incident hashes and triage recommendations to eliminate redundancy. We store any IncidentHash and triage recommendation pairs from the current batch, including those without a triage grade, provided they do not exceed five stored embeddings—aligning with our policy of recommending no more than five similar incidents at a time. New incident embeddings are then saved to the ADLS table for use in the inference pipeline. This procedure repeats until we have accumulated 180 days of historical incident embedding telemetry.
@@ 7. Experiments
We detail our experimental setup and provide a comprehensive analysis of Copilot Guided Response’s performance across each of the three tasks: (1) triaging, where we assess our model’s capability to predict whether an incident is benign, malicious, or informational; (2) investigation, where we evaluate the relevance of our similar incident recommendations in collaboration with security experts; and (3) remediation, where we evaluate our model’s effectiveness in predicting appropriate actions for mitigating security threats.
@@ 7.1. Setup
We present the results for triage and remediation models across a sample of 12 regions, where each regional dataset is divided into three stratified subsets: training, validation, and testing, with splits of 70%, 10%, and 20%, respectively. We perform a grid search to optimize key model parameters on the validation sets as detailed in Section 4.2. While we evaluate each model on its macro-F1 score, we also report precision and recall, as is typical for imbalanced datasets (Freitas et al., 2022, 2021; Duggal et al., 2021, 2020).
Table 2 outlines key statistics for model training and performance across a subset of regions over a two week period. These statistics include a representative sample of the number of unique custom and built-in DetectorIds (rules), the volume of alerts generated by these detectors, and the number of incidents, which are aggregations of alerts. The data reveals significant regional variations, with some regions containing up to 31k unique detector rules across tens of thousands of organizations, substantially complicating the model training process. Furthermore, the distribution of incident sizes, depicted in Figure 2, exhibits a long-tailed pattern where most incidents consist of only a few alerts. However, the larger, more complex incidents pose a greater challenge to triage and identify similar incident recommendations due to their infrequency.
@@ 7.2. GUIDE Dataset
GUIDE encompasses over 13M pieces of evidence across 33 entity types, covering 1.6M alerts and 1M incidents annotated with triage labels from customers over a two week period. Of these incidents, 26k also contain remediation action labels from customers. The dataset includes telemetry across 6.1k organizations, featuring 9.1k unique custom and built-in DetectorIds across numerous security products, encompassing 441 MITRE ATT&CK techniques (Strom et al., 2018)—a representative subset of the results shown in Table 2. We divide the dataset into a train set containing 70% of the data and a test set with 30%, both available as CSV files on Kaggle. By releasing this dataset, we aim to spark new technological innovations in GR, where our system and dataset can serve as critical benchmarks. For additional details on GUIDE, please refer to the Appendix.
@@ 7.3. Triage
The telemetry collection period for triage ranges from 7 to 180 days, varying by region due to computational constraints in incident preprocessing. Table 2 shows that the number of SOC-graded incidents (abbreviated as ”Supp” for support) ranges from 9k to 139k with an imbalance among the three triage classes—19% true positive, 35% false positive, and 46% benign positive (informational). While larger regions tend to have more graded incidents, several factors complicate the number of graded incidents reported during training: (1) larger regions with more organizations and detection rules require more memory during OHE, reducing capacity for additional alerts; (2) the frequency and types of grading vary significantly across SOCs and regions, and (3) we limit the number of graded examples per IncidentHash and triage label to 1k.
Offline results. In Table 2, we present the cross-region model triage performance at the point of maximum Macro-F1 score on the precision-recall curve. The results show an average Macro-F1 score of , with a precision of , and a recall of . This performance demonstrates the model’s ability to effectively manage the complexities of the triage task, ranging from incidents involving a single alert to those comprising hundreds of alerts across numerous detection rules and security products. The vast majority of misclassifications (98.7%) occur when categorizing incidents as BP when they are actually TP or FP. This discrepancy is largely due to the lack of a uniform definition for BP incidents, leading to inconsistent classifications across and within SOCs. However, the more critical misclassifications between TP and FP incidents are relatively rare, accounting for only 1.3% of errors.
@@ 7.4. Investigation
In collaboration with security experts at Microsoft, we manually evaluated our similar incident recommendations due to an absence of a definitive ground truth. For this evaluation, we randomly selected 1k incidents varying in size, detector rules, products, organizations, and regions. Security researchers were then tasked with randomly selecting a similar incident recommendation for each reference incident, and judging its relevance based on criteria such as shared attack patterns, indicators of compromise, and entity types.
Offline results. The assessment found that 94% of the recommended incidents were relevant, with only 2% deemed dissimilar. The quality of recommendations tends to diminish for smaller organizations with limited historical data. Similarly, larger incidents characterized by hundreds of alerts across multiple products and detection rules also show decreased recommendation quality due to their rarity. To address this, a cosine similarity threshold of was identified as the cut-off threshold to ensure that only relevant recommendations are presented to customers. Across regions, we find that 98% of all incidents have one or more recommendations.
@@ 7.5. Remediation
Offline results. We achieve an impressive average cross-region Macro-F1 score of . This high score reflects the relative simplicity of predicting the appropriate remediation action for a single alert compared to the complexities involved in incident triaging.
@@ 8. Deployment
Copilot Guided Response has been successfully deployed across the world, serving thousands of Microsoft Defender XDR customers since its general availability launch in April 2024. To date, CGR has generated millions of guided response recommendations for triage, investigation, and remediation tasks, receiving a ‘positive’ user response rate of 89%, based on the confirmation or dismissal of recommendations.
@@ 9. Conclusion
Copilot Guided Response (CGR) represents the first time a cybersecurity company has openly discussed an industry-scale guided response framework. CGR significantly enhances SOC operations by guiding security analysts through crucial investigation, triaging, and remediation tasks, adeptly handling everything from simple alerts to complex incidents. The performance of CGR has been rigorously evaluated through internal testing, collaboration with Microsoft security experts, and extensive customer feedback, demonstrating its exceptional effectiveness across all three tasks. Deployed globally within Microsoft Defender XDR, CGR generates millions of guided response recommendations weekly, with 89% of user interactions receiving positive feedback. We are optimistic that this research will not only advance guided response capabilities but also encourage greater transparency within the research community.
@@ Acknowledgements.
We thank all of our colleagues who supported this research, including Shira Shacham, Yuval Derman, Oren Saban, Noa Bratman, Inbar Rotem, Omri Kantor, Itamar Karavani, Mari Mishel, Yuval Katav, Niv Zohar, Lior Camri, Leeron Luzzatto, Anna Karp, Ido Nitzan, Corina Feurstein, Ya’ara Cohen, Nadia Tkach Mendes, Pawel Partyka, Shachaf Levy, Blake Strom, along with countless others.References

### v4

@@ 5.2. Triage Recommendations
We leverage the incident embeddings produced during preprocessing, along with the latest version of the triage model, to generate triage recommendations. Each incident is evaluated by the model and given a prediction of true positive (TP), false positive (FP), or benign positive (BP), the latter considered as an informational incident. The confidence of each recommendation is assessed against a precision threshold of to ensure that only reliable recommendations are sent to SOC analysts.
@@ 5.3. Investigation Recommendations
We utilize the incident embeddings produced during the preprocessing step to generate recommendations for similar incidents. This process begins with the retrieval of historical incident embeddings from our Azure Data Lake Storage (ADLS), going back up to 180 days. These embeddings capture past incidents in a vectorized format, enabling efficient comparison. The core of our approach is to match new incidents with historically relevant incidents within the same organization through a three-step matching process:
@@
Top-k similar incident selection. We select the top-k most similar incidents, up to a maximum of five. Exact and cosine similarity matches are ordered, with a higher priority given to exact matches to ensure the most germane comparisons, and ties going to the most recent incident.
@@ 5.4. Remediation Recommendations
Using alert embeddings from preprocessing and the latest remediation model, we generate targeted response actions—contain user, isolate machine, or stop virtual machine—for each alert with confidence above a precision threshold. The system identifies entities (e.g., users, devices, VMs) using encoded rules based on security domain knowledge, and then aggregates the individual alert recommendations into comprehensive incident recommendations.
@@ 6. Embedding Pipeline
We generate historical incident embeddings that allow the similar incident recommendation algorithm to leverage up to 180 days of historical data when making recommendations. Due to the limitations of the training pipeline in processing huge volumes of incident telemetry across regions, we developed a specialized mechanism to generate historical embeddings. This ensures that our similar incident recommendation algorithm rapidly reaches comprehensive historical coverage each time the inference pipeline is executed.
Continuous embedding generation. The embedding pipeline operates in a continuous loop, with each iteration processing data one day further back than the last. Leveraging the preprocessing steps outlined in Section 5.1, we integrate a deduplication process that loads historical incident embeddings and compares incident hashes and triage recommendations to eliminate redundancy. We store any IncidentHash and triage recommendation pairs from the current batch, including those without a triage grade, provided they do not exceed five stored embeddings—aligning with our policy of recommending no more than five similar incidents at a time. New incident embeddings are then saved to the ADLS table for use in the inference pipeline. This procedure repeats until we have 180 days of historical incident embedding telemetry.
@@ 7. Experiments
We evaluate CGR’s performance across three tasks: (1) triaging, assessing the model’s ability to classify incidents as benign, malicious, or informational; (2) investigation, analyzing the relevance of similar incident recommendations; and (3) remediation, evaluating its ability to predict effective threat mitigation actions.
@@ 7.1. Setup
We present results for the triage and remediation models across a sample of 12 regions, where each regional dataset is divided into three stratified subsets: training (70%), validation (10%), and testing (20%). Each run of model training job includes two key parameter optimizations, performed independently for incident triage and alert remediation embeddings:
PCA component selection. We retain the top-k principal components capturing 95% of the data variance, with ranging up to . While the optimal varies across runs and regions, we find that performs well across most scenarios.
Random forest parameter tuning. We optimize key random forest parameters using a grid search (see Section 4.2) We select the model with the highest macro-F1 score on the validation set and report precision and recall metrics, as is standard for imbalanced datasets (Freitas et al., 2022, 2021; Duggal et al., 2021, 2020).
Table 2 summarizes model training and performance statistics over two weeks across select regions, including the count of unique DetectorIds, and alert/incident volumes. Regional variations are significant, with up to 31k detectors across thousands of organizations, complicating the training process. Incident size distribution (Figure 2) is long-tailed, with most incidents comprised of a few alerts, with larger incidents representing a greater challenge for triage and similar incident recommendations due to their rarity.
Limitations. This work focuses on presenting a scalable, unified framework for incident triaging, remediation, and similar incident recommendations, rather than evaluating alternative models or competing security products, which are often undisclosed. The release of the GUIDE dataset enables researchers to explore and optimize new architectures for maximum performance. In addition, CGR is limited to providing recommendations for existing detectors and does not address zero-day or other unmonitored attack vectors.
@@ 7.2. GUIDE Dataset
GUIDE includes over 13M pieces of evidence across 33 entity types, encompassing 1.6M alerts and 1M incidents over a two-week period. Each incident is annotated with a ground-truth triage label by customer security analysts, along with 26k alerts containing labels of remediation actions taken by customers. The dataset is derived from Region 2 (see Table 2) and includes telemetry across 6.1k organizations, featuring 9.1k unique custom and built-in DetectorIds across numerous security products, covering 441 MITRE ATT&CK techniques (Strom et al., 2018). We divide the dataset into a train (70%) and test set (30%), both available as CSV files on Kaggle. The release of GUIDE provides an unparalleled opportunity for the development of GR systems and beyond, with CGR providing a foundational baseline. Additional details on GUIDE can be found in the Appendix.
@@ 7.3. Triage
The telemetry collection period for triage ranges from 7 to 180 days, varying by region due to computational constraints in incident preprocessing. Table 2 shows that the number of SOC-graded incidents (abbreviated as ”Supp” for support) ranges from 8.9k to 139k with an imbalance among the three triage classes of 19% true positive, 35% false positive, and 46% benign positive (informational). While larger regions tend to have more graded incidents, several factors complicate the number of graded incidents reported during training: (1) larger regions with more organizations and detectors require more memory during OHE, reducing capacity for additional alerts; (2) the frequency and types of grading vary significantly across SOCs and regions, and (3) we limit the number of graded examples per IncidentHash and triage label to 1k.
Offline results. In Table 2, we present the cross-region model triage performance at the point of maximum macro-F1 score on the precision-recall curve. The results show an average macro-F1 score of , with a precision of , and a recall of . This performance demonstrates the model’s ability to effectively manage the complexities of the triage task, ranging from incidents involving a single alert to those comprising hundreds of alerts across numerous detectors and security products.
To evaluate CGR’s effectiveness at a more granular level, we examine performance across individual triage classifications in Region 2. The precision-recall curves in Figure 3 show that CGR consistently achieves high AUC scores across triage labels. The majority of misclassifications (81.1%) occur when incidents are categorized as BP instead of TP or FP, or vice versa—a discrepancy mainly due to the lack of a standardized definition for BP incidents, leading to classification inconsistencies across and within SOCs. However, the more critical incident misclassification of TP as BP or FP is rare (2.4%), as shown in Table 3.
@@ 7.4. Investigation
In collaboration with security experts at Microsoft, we manually evaluated our similar incident recommendations due to an absence of a definitive ground truth. For this evaluation, we randomly selected 1k incidents varying in size, detectors, products, organizations, and regions. Security researchers were then tasked with randomly selecting a similar incident recommendation for each reference incident, and judging its relevance based on criteria such as shared attack patterns, indicators of compromise, and entity types.
Offline results. The assessment found that 94% of the recommended incidents were relevant, with only 2% deemed dissimilar. The quality of recommendations tends to diminish for smaller organizations with limited historical data. Similarly, larger incidents characterized by hundreds of alerts across multiple products and detectors also show decreased recommendation quality due to their rarity. To address this, a cosine similarity threshold of was identified as the cut-off threshold to ensure that only relevant recommendations are presented to customers. Across regions, we find that 98% of all incidents have one or more recommendations.
@@ 7.5. Remediation
Offline results. We achieve an impressive average cross-region macro-F1 score of . This high score reflects the relative simplicity of predicting the appropriate remediation action for a single alert compared to the complexities of incident triaging in our dataset.
@@ 8. Deployment
Copilot Guided Response has been successfully deployed across the world, serving thousands of Microsoft Defender XDR customers since its launch in April 2024. CGR has generated millions of guided response recommendations for triage, investigation, and remediation tasks, receiving a ‘positive’ user response rate of 89%, based on the confirmation or dismissal of recommendations.
@@ 9. Conclusion
Copilot Guided Response (CGR) represents the first time a cybersecurity company has openly discussed an industry-scale guided response framework. CGR significantly enhances SOC operations by guiding security analysts through crucial investigation, triaging, and remediation tasks, adeptly handling everything from simple alerts to complex incidents. The performance of CGR has been rigorously evaluated through internal testing, collaboration with Microsoft security experts, and extensive customer feedback, demonstrating its effectiveness across all three tasks. Deployed globally within Microsoft Defender XDR, CGR generates millions of guided response recommendations weekly, with 89% of user interactions receiving positive feedback. In addition, we release GUIDE, the largest publicly available collection of real-world security incidents, comprising 13 million pieces of evidence across one million incidents, each annotated with ground-truth triage labels by customer security analysts. As the first resource of its kind, GUIDE sets a new standard for advancing the development and evaluation of guided response systems and beyond.
@@ Acknowledgements.
We thank our colleagues who supported this research, including Shira Shacham, Yuval Derman, Oren Saban, Noa Bratman, Inbar Rotem, Omri Kantor, Itamar Karavani, Mari Mishel, Yuval Katav, Niv Zohar, Lior Camri, Leeron Luzzatto, Anna Karp, Ido Nitzan, Corina Feurstein, Ya’ara Cohen, Nadia Tkach Mendes, Gil Shmaya, Pawel Partyka, Shachaf Levy, Blake Strom, along with many others.References
@@
- Oliver et al. (2024) Jonathan Oliver, Raghav Batta, Adam Bates, Muhammad Adil Inam, Shelly Mehta, and Shugao Xia. 2024. Carbon Filter: Real-time Alert Triage Using Large Scale Clustering and Fast Search. arXiv preprint arXiv:2405.04691 (2024).